Diffusion Morning Report 05/19/05

Ted Berg tberg at snf.stanford.edu
Thu May 19 06:51:10 PDT 2005


Good Morning All,
    Looking ok this morning. Poly is down -broken tube-replacing. 
Nitride is still under observation for particles. Ebeam still in High vac.
source controller for Tylan 2 is not cooling-investigating.  Everything 
else is OK for now.

  	Date: 	5/19/2005 	THICKNESS-Change at
Tube Name/Number 	Status 	Comments 	 
Tylan 1 	Up 	  	 
Tylan 2 	yellow 	source controller issues 	 
Tylan 3 	Up 	  	 
Tylan 4 	Up 	  	 
Tylan 5-BBr3 	Up 	  	 
Tylan 6-POCl3 	Up 	  	 
Tylan 7-TEOS2 	Up 	  	 
Tylan 8-SiGe 	Up 	  	3.12/30
Tylan 9-Poly 	Down 	tube broken 	6.69/30
Tylan 10-Nitride 	yellow 	Observing particles 	11.86/30
Tylan 11-BPSG-LTO 	Up 	  	18.5/30
Tylan 12-FGA 	Up 	  	 
Tystar 	Up 	  	24.78/30+
Doped Poly 	GONE 	  	 
TEOS 	yellow 	  	 
Tylan 16 	N/A 	  	 
Thermco 1 	Up 	  	 
Thermco 2 	Up 	  	 
Thermco Poly 	yellow 	testing doped 	 
Thermco Nitride 	Down 	Waiting 	 
  	  	  	 
Burn Box 	Up 	  	 
ZYGO 	up 	  	.
  	  	  	 
GAAS 22 	UP 	  	 
GAAS 23 	UP 	  	 
  	  	  	 
SiH4 for furnaces 	  	B on line 400 PSI 	 changeover working well 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20050519/b7443398/attachment.html>


More information about the diffusion mailing list